In studying animal populations we often employ "capturing, marking, releasing, recapturing", using the equation N/M=n/R, where N= total population; M=# of individuals marked; n=# of individuals re-captured without mark; R=# of individuals re-captured with mark. here we are studying a population of rabbits species on an island.
First, estimate the total population of the rabbit species with M=100, n=95, & R=5.
Second, after one year estimate the total population again of the rabbit species with M=200, n=190, & R=10.

Please show all work?
Answer #1

You have to use formula given in the question.

N/M = n/R

1. N/100 = 95/5

N = 19 × 100 = 1900

Total population size is 1900.

2. N/200 = 190/10

N = 19 × 200 = 3800

Total population size after 1 year is 3800.

Please rate high.